OpenStudy (anonymous):

The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Write two equations which represent the two sentences. The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Write two equations which represent the two sentences. @Mathematics

OpenStudy (across):

\[2x+y=46\]\[y=4+x\]

myininaya (myininaya):

P=2x+y y=x+4

myininaya (myininaya):

where P is 46

OpenStudy (across):

meow18, do you know how to solve this system?